# Pros and cons of structural chirality measurements and computation of handedness in periodic solids

### v2

@@ Acknowledgements.
F.G.-O., M.F. and E.B. acknowledge the Fonds de la Recherche Scientifique (FNRS) for financial support, the PDR project CHRYSALID No.40003544 and the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the F.R.S.-FNRS under Grant No. 2.5020.11 and the Tier-1 Lucia supercomputer of the Walloon Region, infrastructure funded by the Walloon Region under the grant agreement No. 1910247. F.G.-O. and E. B. also acknowledge support by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 964931 (TSAR). The work at West Virginia University was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under Award DE‐SC0021375. This work used Bridges2 and Expanse at the Pittsburgh Supercomputer and the San Diego Supercomputer Center through allocation DMR140031 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which National Science Foundation supports grants 2138259, 2138286, 2138307, 2137603, and 2138296.
@@ Appendix A Reciprocal space formulation of the Helicity
This short section is devoted to the straightforward demonstration of Eq. (10). First, it is clear that the value at $\Gamma$ of the Fourier transform of the helicity density operator corresponds to the helicity.
@@
[×$\mathcal{F}[\nabla\times\vec{v}]=i\vec{k}\times\mathcal{F}\left[\vec{v}\right],$ (13)
which completes the prove of Eq. (10).

### v3

@@ Acknowledgements.
F.G.-O., M.F. and E.B. acknowledge the Fonds de la Recherche Scientifique (FNRS) for financial support, the PDR project CHRYSALID No.40003544 and the Consortium des Équipements de Calcul Intensif (CÉCI), funded by the F.R.S.-FNRS under Grant No. 2.5020.11 and the Tier-1 Lucia supercomputer of the Walloon Region, infrastructure funded by the Walloon Region under the grant agreement No. 1910247. F.G.-O. and E. B. also acknowledge support by the European Union’s Horizon 2020 research and innovation program under Grant Agreement No. 964931 (TSAR). F.G.O. also acknowledges financial support from MSCA-PF 101148906 funded by the European Union and the Fonds de la Recherche Scientifique (FNRS) through the grant FNRS-CR 1.B.227.25F The work at West Virginia University was supported by the U.S. Department of Energy (DOE), Office of Science, Basic Energy Sciences (BES), under Award DE‐SC0021375. This work used Bridges2 and Expanse at the Pittsburgh Supercomputer and the San Diego Supercomputer Center through allocation DMR140031 from the Advanced Cyberinfrastructure Coordination Ecosystem: Services & Support (ACCESS) program, which National Science Foundation supports grants 2138259, 2138286, 2138307, 2137603, and 2138296.
@@ Appendix A Reciprocal space formulation of the Helicity
This brief section explores a potential alternative formulation of the helicity operator in reciprocal space.
Notably, the value of the Fourier transform of the helicity density operator at the $\Gamma$ point directly corresponds to the helicity
@@
$\mathcal{F}[\nabla\times\vec{v}]=i\vec{k}\times\mathcal{F}\left[\vec{v}\right],$ (10)
which offers a concise expression for the reciprocal-space formulation of the operator discussed in the main text.